# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ 5.3.1 Model extraction attacks
Such attacks [12, 13, 14] pose a significant threat to DNNs as they enable an adversary to steal the functionality of a victim model. In these attacks, $\mathcal{A}$ queries the victim model (i.e., $\mathcal{M}_{wm}$ in our case) using synthesized samples and collect responses to train a surrogate model (i.e., $\mathcal{M}_{adv}$) to steal $\mathcal{M}_{wm}$’s functionality. The literature on model extraction attacks is limited in the context of LP tasks on GNNs. Therefore, to evaluate GENIE against model extraction attacks, we modify the loss function employed in the knowledge distillation process [61] as outlined in Eq. 11 and Eq. 12.
@@
We consider 3 types of model extraction techniques, viz., soft label, hard label, and double extraction. In soft label extraction, we apply $\mathcal{L}_{CE}$ between the logits of $\mathcal{M}_{wm}$ and $\mathcal{M}_{adv}$ to train $\mathcal{M}_{adv}$ (cf. Eq. 11). In hard label extraction, we apply $\mathcal{L}_{CE}$ between the predictions of $\mathcal{M}_{wm}$ and $\mathcal{M}_{adv}$ to train $\mathcal{M}_{adv}$ (cf. Eq. 12). In double extraction, we perform the hard label extraction twice to obtain the final $\mathcal{M}_{adv}$. Double extraction is a tougher setting since it is difficult for the watermark to survive model extraction twice. We train $\mathcal{M}_{adv}$ model using half of $\mathcal{D}_{test}$ and evaluate it with the other half. Table V shows $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before model extraction and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after model extraction attack.
TABLE V: Impact of model extraction.
@@
We observe a maximum drop of 18.19% from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (for the C.ele dataset) under hard extraction attack, and a maximum drop of 9.82% (for the arXiv dataset) under soft extraction attack. Despite these drops, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains significantly above the watermark threshold (cf. TABLE III) in both cases, which ensures reliable ownership verification. We note a greater drop from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ under hard extraction attack compared to soft extraction attack, which is naturally expected since logits provide richer information about the decision boundary than hard predictions. Despite a more significant drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ and from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ under double extraction attack, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ still remains substantially above the watermark threshold (cf. TABLE III) across all datasets. It demonstrates GENIE’s robustness against persistent model extraction attempts.
Different model architectures: It is also possible that $\mathcal{A}$ might not choose the same architecture to steal the model via model extraction. TABLE VI presents the outcomes of all 3 model extraction attacks when the $\mathcal{M}_{adv}$’s architecture (i.e., GraphSAGE) differs from $\mathcal{M}_{wm}$’s architecture (i.e., GCN). We find that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ is still above the watermark threshold (cf. TABLE III) under all attacks across all datasets; except for the power dataset under double extraction attack. However, $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ in that case drops from 98.93% to 58.81%, rendering the attack useless. Thus, we can conclude that GENIE is robust against model extraction attacks even when $\mathcal{M}_{adv}$ and $\mathcal{M}_{wm}$ have different architecture.
TABLE VI: Impact of model extraction when architecture of $\mathcal{M}_{adv}$ (i.e., GraphSAGE) differs from $\mathcal{M}_{wm}$ (i.e., GCN).
@@ 5.3.2 Knowledge distillation
To test GENIE’s robustness, we assume that the student model have the same architecture as the teacher model. Moreover, we use half of $\mathcal{D}_{test}$ for distillation and evaluate $\mathcal{M}_{adv}$ on the other half. TABLE VII presents $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before knowledge distillation and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after knowledge distillation.
TABLE VII: Impact of knowledge distillation.
@@
We observe a maximum drop of 25.24% from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (for the arXiv dataset). It important to note that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ still remains significantly above the watermark threshold (cf. TABLE III), which indicates that $\mathcal{A}$ was not successful in removing the watermark using knowledge distillation. To summarize, our results show that knowledge distillation was able to transfer the core functionality of the victim model, but the watermark was transferred too (as $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ is still above the threshold for all datasets). We can conclude that GENIE is robust against knowledge distillation.
@@ 5.3.3 Model fine-tuning
FTLL is considered the weakest attack because it has the least capacity to modify the core GNN layers responsible for learning the watermark. Conversely, RTAL is considered the toughest attack because it enables complete fine-tuning of all model layers, providing the highest flexibility to potentially overwrite or distort the watermark embedded across multiple layers. TABLE VIII lists $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before fine-tuning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after fine-tuning with all the four types.
TABLE VIII: Impact of model fine-tuning.
@@ 5.3.4 Model compression
Model pruning: Model or parameter pruning [63] selects a fraction of weights that have the smallest absolute value and makes them zero. It is a computationally inexpensive backdoor defense. We evaluate GENIE against different pruning fractions starting from 0.2 at a step size of 0.2. TABLE IX lists $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before model pruning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after model pruning with different prune percentage. We see that even after pruning 80% [^7] A model obtained after 100% pruning is equal to a random classifier. of $\mathcal{M}_{wm}$’s weights, we are still able to verify the ownership from the resultant $\mathcal{M}_{adv}$ in almost every case. Given that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains above the watermark threshold (cf. TABLE III) roughly in all cases, we conclude that GENIE is robust against model pruning.
TABLE IX: Impact of model pruning.
@@
Weight quantization: It is another model compression technique to reduce the size of a model. It changes the representation of weights to a lower-bit system, thereby saving memory. It is often used to compress large models, e.g., LLMs [64]. We follow the standard weight quantization method [65] with bit-size $=3$. TABLE X reports that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains remarkably above the watermark threshold (cf. TABLE III) after quantization in all the cases, making GENIE is robust against weight quantization.
TABLE X: Impact of weight quantization.
@@ 5.3.5 Fine-pruning
It is a key defense against a backdoor attack that combines model pruning and fine-tuning. It is more effective than individual pruning or fine-tuning, which makes it difficult for the watermark to survive. We start by pruning a fraction of the smallest absolute weights. Next, we fine-tune the pruned model with half of $\mathcal{D}_{test}$ and evaluate the pruned+fine-tuned model (i.e., $\mathcal{M}_{adv}$) with the other half. We perform an exhaustive evaluation with pruning fractions ranging from 0.2-0.8 at a step size of 0.2, which is followed by one of the four types of model fine-tuning (i.e., FTLL, RTLL, FTAL, RTAL). Our rigorous experiments aim to provide a holistic understanding of GENIE’s robustness against the fine-pruning technique. TABLEs XI-XIV exhibit $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before fine-pruning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after fine-pruning with different pruning fractions and fine-tuning methods.
TABLE XI: Impact of pruning + FTLL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 93.75 90.62 81.25 79.68
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 91.21 90.42 87.50 69.82
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 99.77 98.88 97.55 93.11
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 91.71 90.53 85.20 92.30
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 99.00 95.00 87.00 72.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 86.51 81.81 59.24 30.07
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 96.32 88.52 79.43 51.97
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
TABLE XII: Impact of pruning + RTLL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 59.37 57.81 68.75 84.37
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 82.12 82.81 79.58 61.62
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 96.66 97.55 96.22 94.44
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 88.16 88.16 77.51 84.61
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 99.00 93.00 91.99 71.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 44.95 47.04 48.31 39.69
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 59.41 51.42 53.16 46.28
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
We observe the highest drop from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ when fine-pruning is performed using RTAL (cf. TABLE XIV), which is expected as RTAL represents the strongest attack that enables complete fine-tuning of all model layers. Nevertheless, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains above the watermark threshold (cf. TABLE III) in most cases. We strongly believe that GENIE is robust against fine-pruning given that it is failing at only 2 out of 112 cases.
TABLE XIII: Impact of pruning + FTAL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 71.87 60.93 78.12 62.50
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 81.49 73.19 76.31 63.28
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 89.11 84.66 88.66 92.22
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 100 100 99.40 81.65
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 88.00 87.00 84.00 67.99
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 54.45 48.18 33.05 17.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 78.60 75.39 66.20 45.36
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
TABLE XIV: Impact of pruning + RTAL.

### v3

@@ 5.4.1 Model extraction attacks
Such attacks [1, 2, 3] pose a significant threat to DNNs as they enable an adversary to steal the functionality of a victim model. In these attacks, $\mathcal{A}$ queries the victim model (i.e., $\mathcal{M}_{wm}$ in our case) using publicly available test samples and collect responses to train a surrogate model (i.e., $\mathcal{M}_{adv}$) to steal $\mathcal{M}_{wm}$’s functionality. The literature on model extraction attacks is limited in the context of LP tasks on GNNs. Therefore, to evaluate Genie against model extraction attacks, we modify the loss function employed in the knowledge distillation process [59] as outlined in Eq. 6 and Eq. 7.
@@
We consider 3 types of model extraction techniques, viz., soft label, hard label, and double extraction. In soft label extraction, we apply $\mathcal{L}_{CE}$ between the logits of $\mathcal{M}_{wm}$ and $\mathcal{M}_{adv}$ to train $\mathcal{M}_{adv}$ (cf. Eq. 6). In hard label extraction, we apply $\mathcal{L}_{CE}$ between the predictions of $\mathcal{M}_{wm}$ and $\mathcal{M}_{adv}$ to train $\mathcal{M}_{adv}$ (cf. Eq. 7). In double extraction, we perform the hard label extraction twice to obtain the final $\mathcal{M}_{adv}$. Double extraction is a tougher setting since it is difficult for the watermark to survive model extraction twice. We train $\mathcal{M}_{adv}$ model using half of $\mathcal{D}_{test}$ and evaluate it with the other half. Table VI shows $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before model extraction and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after model extraction attack.
TABLE VI: Impact of model extraction.
@@
We see a maximum drop of 18.19% from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (for the C.ele dataset) under hard extraction attack, and a maximum drop of 9.82% (for the arXiv dataset) under soft extraction attack. Despite these drops, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains significantly above the watermark threshold (cf. TABLE III) in both cases, which ensures reliable ownership verification. We note a greater drop from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ under hard extraction attack compared to soft extraction attack, which is naturally expected since logits provide richer information about the decision boundary than hard predictions. We observe a more significant drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ and from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ under double extraction attack. The drop in performance can be anticipated due to the fact that the first iteration of model extraction creates a data drift from its initial distribution [60]. This drift then introduces an error that propagates in the second iteration, causing an expected decrease in performance [9]. However $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains substantially above the threshold for all attacks demonstrating Genie’s robustness against persistent model extraction attempts.
Different model architectures: It is also possible that $\mathcal{A}$ might not choose the same architecture to steal the model via model extraction. TABLE VII presents the outcomes of all 3 model extraction attacks when the $\mathcal{M}_{adv}$’s architecture (i.e., GraphSAGE) differs from $\mathcal{M}_{wm}$’s architecture (i.e., GCN). We find that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ is still above the watermark threshold (cf. TABLE III) under all attacks across all datasets; except for the power dataset under double extraction attack. However, $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ in that case drops from 98.93% to 58.81%, rendering the attack useless. Thus, we can conclude that Genie is robust against model extraction attacks even when $\mathcal{M}_{adv}$ and $\mathcal{M}_{wm}$ have different architecture.
TABLE VII: Impact of model extraction when architecture of $\mathcal{M}_{adv}$ (i.e., GraphSAGE) differs from $\mathcal{M}_{wm}$ (i.e., GCN).
@@ 5.4.2 Knowledge distillation
To test Genie’s robustness, we assume that the student model have the same architecture as the teacher model. Moreover, we use half of $\mathcal{D}_{test}$ for distillation and evaluate $\mathcal{M}_{adv}$ on the other half. TABLE VIII presents $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before knowledge distillation and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after knowledge distillation.
TABLE VIII: Impact of knowledge distillation.
@@
We see a maximum drop of 25.24% from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (for the arXiv dataset). It important to note that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ still remains significantly above the watermark threshold (cf. TABLE III), which indicates that $\mathcal{A}$ was not successful in removing the watermark using knowledge distillation. To summarize, our results show that knowledge distillation was able to transfer the core functionality of the victim model, but the watermark was transferred too (as $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ is still above the threshold for all datasets). We can conclude that Genie is robust against knowledge distillation.
@@ 5.4.3 Model fine-tuning
FTLL is considered the weakest attack because it has the least capacity to modify the core GNN layers responsible for learning the watermark. Conversely, RTAL is considered the toughest attack because it enables complete fine-tuning of all model layers, providing the highest flexibility to potentially overwrite or distort the watermark embedded across multiple layers. TABLE IX lists $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before fine-tuning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after fine-tuning with all the four types.
TABLE IX: Impact of model fine-tuning.
@@ 5.4.4 Model compression
Model pruning: Model or parameter pruning [62] selects a fraction of weights that have the smallest absolute value and makes them zero. It is a computationally inexpensive watermark removal technique. We evaluate Genie against different pruning fractions starting from 0.2 at a step size of 0.2. TABLE X lists $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before model pruning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after model pruning with different prune percentage. We see that even after pruning 80% [^6] A model obtained after 100% pruning is equal to a random classifier. of $\mathcal{M}_{wm}$’s weights, we are still able to verify the ownership from the resultant $\mathcal{M}_{adv}$ in almost every case. Given that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains above the watermark threshold (cf. TABLE III) roughly in all cases, we conclude that Genie is robust against model pruning.
TABLE X: Impact of model pruning.
@@
Weight quantization: It is another model compression technique to reduce the size of a model. It changes the representation of weights to a lower-bit system, thereby saving memory. It is often used to compress large models, e.g., LLMs [63]. We follow the standard weight quantization method [64] with bit-size $=3$. TABLE XI reports that $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains remarkably above the watermark threshold (cf. TABLE III) after quantization in all the cases, making Genie is robust against weight quantization.
TABLE XI: Impact of weight quantization.
@@ 5.4.5 Fine-pruning
It is a key defense against a backdoor attack that combines model pruning and fine-tuning. It is more effective than individual pruning or fine-tuning, which makes it difficult for the watermark to survive. We start by pruning a fraction of the smallest absolute weights. Next, we fine-tune the pruned model with half of $\mathcal{D}_{test}$ and evaluate the pruned+fine-tuned model (i.e., $\mathcal{M}_{adv}$) with the other half. We perform an exhaustive evaluation with pruning fractions ranging from 0.2-0.8 at a step size of 0.2, which is followed by one of the four types of model fine-tuning (i.e., FTLL, RTLL, FTAL, RTAL). Our rigorous experiments aim to provide a holistic understanding of Genie’s robustness against the fine-pruning technique. TABLE XII exhibits $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ before fine-pruning (in column with $\star$) and $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ after fine-pruning with different pruning fractions followed by RTAL fine-tuning method (cf. TABLEs D.1-D.3 in Appendix D.2 for fine-pruning via FTLL, RTLL, and FTAL, respectively).
We observe the highest drop from $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ to $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ when fine-pruning is performed using RTAL (cf. TABLE XII), which is expected as RTAL represents the strongest attack that enables complete fine-tuning of all model layers. Nevertheless, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ remains above the watermark threshold (cf. TABLE III) in most cases. We strongly believe that Genie is robust against fine-pruning given that it is failing at only 2 out of 112 cases.
TABLE XII: Impact of pruning + RTAL.